Suppose the price of pork fell from $3.00 to $2.50 per pound. Over the same period, the CPI (which reflects the price of all consumer goods) increased by 10%. What is the percentage change in price (%) that we should use to calculate the price elasticity of demand for beef? (Hint: Recall that we use the midpoint formula to calculate elasticities.)

Suppose the price of lettuce rises by 20%. As a result, the quantity of lettuce demanded decreases by 5%. What is (the absolute value of) the price elasticity of demand for lettuce?
Answer #1

a) % change in the price of the goods = ( 2.50 - 3 ) / (2.50 + 3) / 2

= 0.50 / 2.75

= 0.18 or we can say that the price of the pork changed by 18%.

b) Price elasticity of the goods = % change in demand / % change in price

= 5 /20 = 0.25 is the price elasticity of lettuce.